OpenStudy (anonymous):

10. If an employee walks at 2/3 of his usual speed, he gets late to the office by 15 mins. What time does he usually take to reach the office? A. 10 min. B. 30 min C. 25 min, D. 35 min

OpenStudy (anonymous):

the distance is the same he walks D D= speed *time= say x*t originally D=(2/3)*x (t+15) equate the distances xt=(2/3)*x)(t+15) let x=6 it could be anything 6t=4(t+15) 2t=60 t=30 the time it normally takes check my reasoning

OpenStudy (anonymous):

thanks for explaining